write the equation of a line has the same slope as y=2x+3 that passes through the point (4,0)
Answer by jim_thompson5910(35256)

y = mx+b

y = 2x+b

0 = 2(4) + b

0 = 8 + b

-8 = b

b = -8


So the equation is y = 2x-8